4. Last year Hamdi Corp. had sales of $500, 000, operating costs of $450,000, and year-e assets of $395,000. The debt-to-total-assets ratio was 178, the interest rate on the debt was 7.58, and the firms tax rate was 35%. The new CFO wants to see how the ROE would have been affected if the firm had used a 50% debt ratio. Assume that sales, operating costs, total assets, and the tax rate would not be affected, but the interest rate would rise to 8.0By how much would the ROE change in response to the change in the capital structure? (18 points)
Answer #1

Answer:

Debt to total asset ratio = 17%

Total assets = $395,000

Debt =395000 * 17% = $67,150

Equity = 395000 - 67150 = $327,850

Total asset remaining same if debt ratio = 50%

Debt = 395000 * 50% = $197,500

Equity =39500 -197500 =$197,500

Last Year Proposed Sales Operating cost EBIT Interest expense $67,150 *7.5% EBT Tax at 35% Net Income (NI $500,000.00 $450,00

As such due to change in capital structure:

Increase in ROE = 11.26% - 8.91% = 2.34%

Increase in ROE as % = 2.34% / 8.91% = 26.26%